# Optimizing Symbol Visibility through Displacement

### v2

@@ 3.2 Computing staircases with gap arbitrarily close to the supremum
We next prove that for every reasonable layout with gap $g$, there is a staircase with a gap at least $g-\delta$, for any $\delta>0$. Moreover, with $\gamma^{\star}$ being the supremum gap over all layouts, a staircase of gap $\gamma^{\star}-\delta$ can be efficiently computed.
For this, we first look at staircases in more detail. Consider a staircase of $n$ squares, facing right and up, with centroids $(x_{i},y_{i})$, $1\leq i\leq n$. We define $\Delta y_{i}=y_{i+1}-y_{i}>0$ and $\Delta x_{i}=x_{i+1}-x_{i}\geq 0$, for $11\leq i\leq n-$ 1. If $\Delta x_{i}>0$, the left and lower sides of $s_{i}$ are fully visible, and the gap of $s_{i}$ is $\Delta y_{i}+\Delta x_{i}$; the top square $s_{n}$ has gap $2$. If all $\Delta x_{i}$ are positive, the staircase is called proper.
Now consider the problem of finding such a proper staircase of large gap. For this, the $\Delta y_{i}$ are fixed, but $x_{1}<x_{2}<\cdots<x_{n}$ can be chosen freely, meaning that the values $\Delta x_{i}$ can be any positive numbers satisfying $\sum_{i=1}^{n-1}\Delta x_{i}\leq w-1$. We want to maximize the gap $\min_{i=1}^{n-1}(\Delta y_{i}+\Delta x_{i})$ subject to the previously mentioned constraints. Due to the strict inequalities on the $\Delta x_{i}$’s, the maximum may not exist (as pointed out in Section 2). But allowing $\Delta x_{i}\geq 0$, the maximum is attained by the solution of a linear program:
@@
Now we are prepared for the main result of this section which also implies that the optimal solution $g^{\star}$ of this linear program equals $\gamma^{\star}$, the supremum gap over all layouts.
@@
Let $\mathcal{A}$ be a reasonable layout with gap $g$. There exists a feasible solution of the linear program (1) with value at least $g$. Moreover, for every $\delta>0$, there exists a proper staircase $\mathcal{A}^{\prime}$ with gap at least $g-\delta$.
Proof 3.7.
@@
The $\Delta Y_{i}$ rectangles are further subdivided into blocks whose intervals are gaps between vertically adjacent squares. For example, $\Delta Y_{4}$ is made of two blocks. The interval of the orange-black block is the vertical gap between the lower sides of $t_{4}$ (orange) and $t_{1}$ (black), while the interval of the black-green block is the vertical gap between the lower sides of $t_{1}$ (black) and $t_{5}$ (green). Thus, each block interval is of the form $\Delta y_{j}=y_{j+1}-y_{j}$.
As argued in the proof of Lemma 3.3, the $\Delta X_{i}$ and $\Delta Y_{i}$ satisfy the constraints of the linear program (1), with $g$ being the gap of $\mathcal{A}$.
Figure 6: The proof of Lemma 3.6.
We will perform discrete steps that gradually turn the $\Delta Y_{i}$ into the prescribed $\Delta y_{i}=y_{i+1}-y_{i}$, while changing the $\Delta X_{i}$ into suitable $\Delta x_{i}$. If we can maintain the constraints of (1) throughout, we will arrive at a feasible solution of (1)—that we can interpret as a staircase—with value at least $g$. From this, we can construct a proper staircase with gap at least $g-\delta$, by slightly redistributing the $\Delta x_{i}$ to make all of them positive.
We point out that the $\Delta X_{i},\Delta Y_{i}$ are sorted by stacking order; our stepwise process will first result in values $\Delta X^{\prime}_{i},\Delta Y^{\prime}_{i}$ such that the $\Delta Y^{\prime}_{i}$ are a permutation of the $\Delta y_{i}$. This means, we still have to sort the values accordingly before we can interpret the solution of (1) as a staircase. As the linear program is agnostic to permutations, this does not change the gap.
If the $\Delta Y_{i}$ are already a permutation of the $\Delta y_{i}$, we are done after sorting them (see the end of the proof below). This is the case if and only if each $\Delta Y_{i}$ consists of exactly one block.
@@
Now we repeatedly move blocks from rectangles with at least two blocks to rectangles with no block. We can visually analyze this as follows: think of a basin that initially holds the $\Delta Y_{i}$ rectangles, $i=1,\ldots,n-1$, as in part (a) of Figure 6. For each $i$, we pour $\Delta X_{i}$ units of water into the basin, which corresponds to the area of the rectangles with colored boundary in part (a) of Figure 6. As $\min_{i}(\Delta X_{i}+\Delta Y_{i})\geq g$, the water will settle at some level $\geq g$; see part (b) of the figure. Moving a block to a “free slot” will submerge it further, and this can only increase the water level; see step (b)-(c).
In the end, we have one block $\Delta y_{j}=y_{j+1}-y_{j}$ per slot, and sorting the slots by index as in part (d) yields rectangles $\Delta y_{1},\ldots,\Delta y_{n-1}$, with columns $\Delta x_{1},\ldots\Delta x_{n-1}$ of water above them, such that $\min(\Delta x_{i}+\Delta y_{i})\geq g$. In general, some $\Delta y_{i}$’s can still be above the water level in which case the corresponding $\Delta x_{i}$ is 00. This is our desired solution of (1) from which we can in turn build a staircase with the prescribed $x$- and $y$-gaps (upper right part of the figure).
@@ 4 Squares stabbed by a vertical line
In $O(n)$ time, we now construct a layout for all squares, of gap roughly $\frac{\delta}{2}$, to prove the statement. To do so, we “squeeze” the layouts in individual buckets appropriately. We assume w.l.o.g. that the even bucket staircases are facing right and up, while the odd ones are facing left and up, as in Figure 8 (left).
Figure 7: Squeezing staircase layouts
Figure 7: Squeezing staircase layouts
Figure 8: Arranging $n$ uniformly spaced squares in a zigzag layout
@@
See Figure 8. We place bundles of $\lfloor k\rfloor$ squares each, as indicated in the figure, starting from the lowest one. This layout uses precisely the $2\lfloor k\rfloor$ $x$-coordinates $\frac{1}{2}+\frac{i}{2\lfloor k\rfloor-1},i=0,\ldots,2\lfloor k\rfloor-1$. This means that every square has $x$-gap at least $\frac{1}{2\lfloor k\rfloor-1}\geq\frac{1}{2k-1}$. The $y$-gap is at least $1/k$ for each square, due to uniform spacing. Both gaps are attained for example by the second-lowest square, so the bound in the lemma cannot be improved for this layout.
@@
The first one is that an upper corner and a lower corner of $s$ are covered. These could be adjacent corners (with some part of the connecting side visible), or antipodal corners as in Figure 11. By Lemma 4.7, one of the adjacent squares must be behind $s$; w.l.o.g. it is the next higher one $b$ (blue).
Figure 10: Case 1: A non-standard bad square (black) with an upper and a lower corner covered
Figure 10: Case 1: A non-standard bad square (black) with an upper and a lower corner covered
Figure 11: Case 2: A non-standard bad square (black) with two upper or two lower corners covered
@@
Figure 12: Counting standard bad squares with centers in a vertical window of height $1$
@@
Figure 13: Proof of bitone stacking order
Refer to captionRefer to captionRefer to caption
Figure 13: Proof of bitone stacking order
@@
Proof 4.17.
A square with at least three visible corners (and only such squares remain) is called down square if the lower side is fully visible, and up square if the upper side is fully visible. A top square is both up and down.
@@
Except $O(\log k)$ of them (the ones directly below a removed bad square), all squares have $y$-gap $1/k$. Let $\varepsilon$ be the minimum $x$-gap among these squares. Then the layout has gap at most $1/k+\varepsilon$. But we know that the sum of all $x$-gaps is at most $1$ (they don’t overlap and “live” outside the top square), so the minimum $x$-gap is $1/(2k-O(\log k))$ which proves the theorem in this case.
The other case is that there are two consecutive squares of different types with centers in the subwindow, see Figure 14 (middle). In this case, we have a generalized staircase. We let $s$ be the lower one and $t$ the higher of the two squares, and we zoom in on the situation; see Figure 14 (right).
Both squares stick out of the ones up to $1$ higher than $s$ (otherwise, they can’t have $3$ visible corners). On the other hand, the squares up to $1$ lower than $t$ stick out of both $s$ and $t$. It follows that—as in the previous case—the $x$-gaps of all involved squares live outside of the top square among them, and they do not overlap (i.e. they are disjoint). More specifically, the $x$-gaps of the squares above $s$ live in the orange regions in Figure 14, while the $x$-gaps of the squares below $t$ live in the blue regions.
@@ 5 Conclusion
We initiated the algorithmic study of optimizing the visibility of overlapping symbols by finding both a suitable drawing order and a limited displacement. This novel setting leads to various interesting and challenging problems. In this paper we focused solely on unit squares, presented structural insights, as well as several intricate approximation algorithms.
@@
It is natural to wonder if the stacking order of every optimal solution follows its $y$-order. However, this is not always the case, refer to Figure 15 for an illustration: the optimal layout (leftmost figure) is better than the best layout when stacking order follows the $y$-order or the inverse $y$-order (second and third figures), which is better than the best layout among all staircases (rightmost figure).
Figure 15: Four layouts for $y=\{0.5,0.7,0.8,1.25,1.35,1.45,1.55,1.65,1.75\}$. The minimum visible perimeter is indicated below each layout, the stacking order by the numbers in the corners of each square. From left to right: optimal layout; optimal layout with a stacking order matching the $y$-order; optimal layout with a stacking order matching the inverse $y$-order; optimal staircase. These layouts were computed via (I)LPs using a difference of $0.001$ to turn strict inequalities into non-strict inequalities.

### v3

@@ 3.2 Computing staircases with gap arbitrarily close to the supremum
We next prove that for every reasonable layout with gap $\gamma$, there is a staircase with a gap at least $\gamma-\delta$, for any $\delta>0$. Moreover, with $\gamma^{\star}$ being the supremum gap over all layouts, a staircase of gap $\gamma^{\star}-\delta$ can be efficiently computed.
For this, we first look at staircases in more detail. Consider a staircase of $n$ squares, facing up and right, with centroids $(x_{i},y_{i})$, $1\leq i\leq n$. We define $\Delta y_{i}=y_{i+1}-y_{i}>0$ and $\Delta x_{i}=x_{i+1}-x_{i}\geq 0$, for $11\leq i\leq n-$ 1. If $\Delta x_{i}>0$, the left and lower sides of $s_{i}$ are fully visible, and the gap of $s_{i}$ is $\Delta y_{i}+\Delta x_{i}$; the top square $s_{n}$ has gap $2$. If all $\Delta x_{i}$ are positive, the staircase is called proper.
Now consider the problem of finding such a proper staircase of large gap. For this, the $\Delta y_{i}$ are fixed, but $x_{1}<x_{2}<\cdots<x_{n}$ can be chosen freely, meaning that the values $\Delta x_{i}$ can be any positive numbers satisfying $\sum_{i=1}^{n-1}\Delta x_{i}\leq w-1$. We want to maximize the gap $\min_{i=1}^{n-1}(\Delta y_{i}+\Delta x_{i})$ subject to the previously mentioned constraints. Due to the strict inequalities on the $\Delta x_{i}$’s, the maximum may not exist (as pointed out in Section 2). But allowing $\Delta x_{i}\geq 0$, the maximum is attained by the solution of a linear program in the variables $\Delta x_{i}$ and $g$ (the $\Delta y_{i}=y_{i+1}-y_{i}$ are constants):
@@
We formalize this in the main result of this section which also implies that the optimal solution $g^{\star}$ of this linear program equals $\gamma^{\star}$, the supremum gap over all layouts.
@@
Let $\mathcal{A}$ be a reasonable layout with gap $\gamma$. There is a feasible solution of the linear program (1) with value $g\geq\gamma$. Vice versa, for every feasible solution of (1) with value $g$, and for every $\delta>0$, there exists a proper staircase $\mathcal{A}^{\prime}$ with gap $g-\delta$.
Proof 3.8.
@@
Figure 6: Proof of Lemma 3.7.
The $\Delta Y_{i}$ rectangles are further subdivided into blocks whose intervals are gaps between vertically adjacent squares. For example, $\Delta Y_{4}$ is made of two blocks. The interval of the orange-black block is the vertical gap between the lower sides of $t_{4}$ (orange) at height $y_{2}-1/2$ and $t_{1}$ (black) at height $y_{3}-1/2$; the interval of the black-green block is the vertical gap between the lower sides of $t_{1}$ (black) at height $y_{3}-1/2$ and $t_{5}$ (green) at height $y_{4}-1/2$. For the green-red block, the interval corresponds to the gap between the upper sides of $t_{5}$ (green) at height $y_{4}+1/2$ and $t_{2}$ (red) at height $y_{5}+1/2$. In general, each block interval is therefore of the form $\Delta y_{j}=y_{j+1}-y_{j}$.
As argued in the proof of Lemma 3.3, the $\Delta X_{i}$ and $\Delta Y_{i}$ satisfy the inequalities
$\Delta X_{i}+\Delta Y_{i}\geq\gamma,\quad i=1,\ldots,n-1,$ (2)
with $\gamma$ being the gap of $\mathcal{A}$.
We will perform discrete steps that gradually turn the $\Delta Y_{i}$ into the prescribed $\Delta y_{i}=y_{i+1}-y_{i}$, while changing the $\Delta X_{i}$ into suitable nonnegative $\Delta x_{i}$, maintaining their sum. If we can also maintain the constraints (2) throughout, we will arrive at a feasible solution of (1) with value $g\geq\gamma$.
Initially, the $\Delta X_{i},\Delta Y_{i}$ are sorted by stacking order; our step-wise process will therefore first result in values $\Delta X^{\prime}_{i},\Delta Y^{\prime}_{i}$ such that the $\Delta Y^{\prime}_{i}$ are a permutation of the $\Delta y_{i}$. Without affecting the constraints (2), we finally reshuffle the pairs $(\Delta X^{\prime}_{i},\Delta Y_{i})$ to get a solution of (1).
Here is how the step-wise process works. If the $\Delta Y_{i}$ are already a permutation of the $\Delta y_{i}$, we are done after reshuffling. This is the case if and only if each $\Delta Y_{i}$ consists of exactly one block.
@@
Now we repeatedly move blocks from rectangles with at least two blocks to rectangles with no block. We can visually analyze this as follows: think of a basin that initially holds the $\Delta Y_{i}$ rectangles, $i=1,\ldots,n-1$, as in part (a) of Figure 6. For each $i$, we pour $\Delta X_{i}$ units of water into the basin, which corresponds to the area of the rectangles with colored boundary in part (a) of Figure 6. As $\bar{\gamma}:=\min_{i}(\Delta X_{i}+\Delta Y_{i})\geq\gamma$, the water will settle at level at least $\bar{\gamma}\geq\gamma$; see part (b) of the figure. Moving a block to a “free slot” will submerge it further, and this can only increase the water level; see step (b)-(c).
In the end, we have one block $\Delta y_{j}=y_{j+1}-y_{j}$ per slot, and sorting the slots by index $j$, as in part (d), yields rectangles $\Delta y_{1},\ldots,\Delta y_{n-1}$, with columns $\Delta x_{1},\ldots\Delta x_{n-1}$ of water above them, such that $\min(\Delta x_{i}+\Delta y_{i})\geq\bar{\gamma}$. In general, some $\Delta y_{i}$’s can still be above the water level in which case the corresponding $\Delta x_{i}$ is 00. This is our desired solution of (1).
For the “vice versa” statement of the lemma, consider any feasible solution of (1) with value $g$. From this, we can construct a proper staircase with gap at least $g-\delta$, by slightly redistributing the $\Delta x_{i}$ to make all of them positive. Then, we can in turn build a staircase with the prescribed $x$- and $y$-gaps as in the upper right part of the figure.
@@ 4 Squares stabbed by a vertical line
In $O(n)$ time, we now construct a layout for all squares, of gap roughly $\frac{\delta}{2}$, to prove the statement. To do so, we “squeeze” the layouts in individual buckets appropriately. We assume w.l.o.g. that the even bucket staircases are facing up and right, while the odd ones are facing up and left, as in Figure 8 (left).
Figure 7: Squeezing staircase layouts.
Figure 8: Arranging $n$ uniformly spaced squares in a zigzag layout.
@@
We place bundles of $\lfloor k\rfloor$ squares each, as indicated in Figure 8, starting from the lowest one. This layout uses precisely the $2\lfloor k\rfloor$ $x$-coordinates $\frac{1}{2}+\frac{i}{2\lfloor k\rfloor-1},i=0,\ldots,2\lfloor k\rfloor-1$. This means that every square has $x$-gap at least $\frac{1}{2\lfloor k\rfloor-1}\geq\frac{1}{2k-1}$. The $y$-gap is at least $1/k$ for each square, due to uniform spacing. Both gaps are attained for example by the second-lowest square, so the bound in the lemma cannot be improved for this layout.
@@
The first one is that an upper corner and a lower corner of $s$ are covered. These could be adjacent corners (with some part of the connecting side visible), or antipodal corners as in Figure 11. In both of those instances, by Lemma 4.7, one of the adjacent squares must be behind $s$; w.l.o.g. it is the next higher one $b$ (blue).
Figure 10: Case 1: A non-standard bad square (black) with an upper and a lower corner covered.
Figure 11: Case 2: A non-standard bad square (black) with two upper or two lower corners covered.
@@
Figure 12: Counting standard bad squares with centers in a vertical window of height $1$.
@@
Proof 4.17.
A square with at least three visible corners (and only such squares remain) is called a down square if the lower side is fully visible, and an up square if the upper side is fully visible. A top square is both up and down.
Figure 13: Proof of the bitone stacking order. The visible perimeter of $s_{i}$ is marked in bold.
@@
Except $O(\log k)$ of them (the ones directly below a removed bad square), all squares have $y$-gap $1/k$. Let $\varepsilon$ be the minimum $x$-gap among these squares. Then the layout has gap at most $1/k+\varepsilon$. But we know that the sum of all $x$-gaps is at most $1$ (they do not overlap and “live” outside the top square), so the minimum $x$-gap is $1/(2k-O(\log k))$ which proves the theorem in this case.
The other case is that there are two consecutive squares of different types with centers in the subwindow, see Figure 14 (middle). In this case, we have a generalized staircase. We let $s_{i}$ be the lower one and $s_{i^{\prime}}$ the higher of the two squares, and we zoom in on the situation; see Figure 14 (right).
Both $s_{i}$ and $s_{i^{\prime}}$ stick out of the squares $s_{j}$, $j>i^{\prime}$, that are up to $1$ higher than $s_{i}$ (otherwise, they cannot have $3$ visible corners). On the other hand, the squares $s_{j}$, $j<i$, that are up to $1$ lower than $s_{i^{\prime}}$ stick out of both $s_{i}$ and $s_{i^{\prime}}$. It follows that—as in the previous case—the $x$-gaps of all involved squares live outside of the top square among them, and they do not overlap (i.e. they are disjoint). More specifically, the $x$-gaps of the squares above $s_{i}$ live in the orange regions in Figure 14, while the $x$-gaps of the squares below $s_{i^{\prime}}$ live in the blue regions.
@@ 5 Conclusion
We initiated the algorithmic study of optimizing the visibility of overlapping symbols by finding both a suitable drawing order and a limited displacement. This novel setting leads to various interesting and challenging problems. In this paper, we focused solely on unit squares, presented structural insights, as well as several intricate approximation algorithms.
@@
It is natural to wonder if the stacking order of every optimal solution follows its $y$-order. However, this is not always the case, refer to Figure 15: the optimal layout (leftmost figure) is better than the best layout when the stacking order follows the $y$-order or the inverse $y$-order (second and third figures), which is better than the best layout among all staircases (rightmost figure).
Figure 15: Four layouts for $y=\{0.5,0.7,0.8,1.25,1.35,1.45,1.55,1.65,1.75\}$ in a strip of width 2. The minimum visible perimeter is indicated below each layout, with the red squares having exactly this visible perimeter. The stacking order is given by the numbers in the corners of each square. From left to right: optimal layout; optimal layout with a stacking order matching the $y$-order; optimal layout with a stacking order matching the inverse $y$-order; optimal staircase. These layouts were computed via (I)LPs using a difference of $0.001$ to turn strict inequalities into non-strict inequalities.
Our results assume a strip of width at most 2, as well as distinct $y$-coordinates. While we can apply our constructions to the more general case, they may be of inferior quality. We leave to future work to establish whether and how these assumptions can be lifted and to determine the computational complexity of the problems in this area.